What is the corresponding annual effective interest rate equivalent to an effective annual discount

rate of 5% p.a.?


1
Expert's answer
2021-04-28T16:24:15-0400

Amount of Rs.100 for 1 year when compounded half-yearly} ;


"=Rs. [100+(1+\frac {3} {100})^{2}] =Rs. 106.09"


∴ Effective rate =(106.09−100)% =6.09%